# Enhancing Quantum Memory Lifetime with Measurement-Free Local Error Correction and Reinforcement Learning

### v1

@@ VI.2 Reducing rate of mid-circuit readouts
Using only LEC in 2D toric code has limitations for a practical application. In 2D toric code, memory lifetime extended by LEC does not scale with $L$ as observed in Fig. 10. Also, undecoded error chains get accumulated by the repetition of LECs. These imply that extending memory lifetime with 2D toric code has some upper limit. However, we can envision a hybrid decoding strategy, combining the advantages of both LEC and global decoding in 2D toric code. In particular, repeated application of LEC can be used to reduce the rate of error accumulation of an encoded logical qubit. Then, we can periodically perform the mid-circuit readouts of ancillas before a round of global decoding. The key advantage of this approach is that LEC circuits can be applied more quickly compared to the global decoding that involves mid-circuit readouts.
@@
In practice, we aim the memory to exhibit a logical error rate $P_{log}$ below some target rate $P_{targ}$ after a given physical time $T$ where each time step introduces some $p_{err}$ to data qubits. We compare original and hybrid decoding protocols as in Fig. 12(a). The original decoding protocol performs global decoding for $N_{MWPM}^{(original)}$ times with equally spaced time intervals throughout the total time $T$. On the other hand, the hybrid decoding protocol performs global decoding for $N_{MWPM}^{(hybrid)}$ with RL-optimized LEC circuits interleaved between global decoding. Note that each global decoding requires $L$ repetitions of mid-circuit readout of ancillas for faulty syndrome extraction before the 3D MWPM algorithm. Thus, our goal is to find the regime of parameters where
$N_{MWPM}^{(original)}>N_{MWPM}^{(hybrid)}.$ (14)
We focus on the system including multi-qubit gates with $p_{gate}=1\times 10^{-4}$, where the fidelity of the CNOT gate is $99.96\%$ and that of the CCX/CCZ gate is $99.94\%$. Note that from our choice of the error model, fidelity of CNOT gate $F_{2Q}=1-4\times p_{gate}$, and fidelity of CCX/CCZ gate $F_{3Q}=1-6\times p_{gate}$. In this regime, we investigate what is the minimum number of global decoding for each decoding protocol to achieve the final logical error rate after total time $T$ to be below $15\%$. As shown in Fig. 12(b), we perform 1D and 2D optimization for the original and hybrid decoding protocol, respectively. For the original decoding, we vary $N_{MWPM}$ and check the final $P_{L}$ after time $T$. On the other hand, for the hybrid decoding, we first fix a total number of global decoding and then find $N_{LEC}$ that minimizes $P_{L}$ after time $T$. Then, we choose the minimum number of global decoding for each of the protocols that satisfy $P_{L}<15\%$.
@@
We obtain the comparison between the two protocols as shown in Fig. 12(c). When our 2D toric code lattice has $L=8$, the hybrid decoding can reduce the number of global decoding for the constant amount regardless of $T$. In other words, the gain of using the hybrid decoding protocol gets decreased as we want to preserve the memory for a longer time. However, for a larger lattice with $L=16$, we observe that the hybrid decoding can reduce the number of global decoding (and thus mid-circuit readouts) by about half regardless of how long we want to preserve the memory. We also compare minimum $N_{MWPM}$ for larger $p_{gate}$ as well. When $p_{gate}=3\times 10^{-3}$, i.e. $F_{2Q}=99.88\%$ and $F_{3Q}=99.82\%$, we obtain that the hybrid protocol can reduce the number of global decoding by 1 or 2 compared to the original protocol regardless of $T$. Note that we observe the same result, even though the original decoding protocol is improved by replacing the 3D MWPM following repeated faulty syndrome extractions with the 2D MWPM following a perfect syndrome extraction. Thus, we can conclude that using an LEC circuit reduces the number of mid-circuit readouts by $L\times(N_{MWPM}^{(original)}-N_{MWPM}^{(hybrid)})$ throughout the entire physical time, i.e. we can replace some portion of the ancilla readouts into possibly more cost-efficient operations. Our investigation confirms that there exists a regime of high multi-qubit gate fidelity where LEC can be practically useful for a memory lifetime enhancement.
VI.3 Dissipatively preparing topological phases
@@ VII Discussion and Outlook
In this work, we develop an RL framework to optimize measurement-free error-correcting circuits, composed of a finite number of faulty multi-qubit gates. We specifically focus on a finite-sized 2D toric code, 2D Ising model, and 4D toric code. By introducing a notion of effective code distance $D_{eff}$, we find that when gate fidelities are high enough, the RL-optimized LEC circuits are better than conventional LEC circuits in extending the lifetime of a memory encoded in these codes. Then, we additionally investigate two potential applications of this LEC scheme. First, we show that the portion of global decoding, which consists of mid-circuit readout and classical decoding, can be replaced with LEC circuits with high-fidelity multi-qubit gates for preserving 2D toric code memory. Also, we discuss that quantum states with topological phases can be prepared dissipatively.
@@
Our RL framework is adaptive to the needs of various improvements or modifications. For example, keeping the same gate operations as an input action space, we can simply substitute a biased, realistic error model into our current noise channel. Since coherent errors turn into Pauli errors when we trace out the ancillas at the ancilla reset step [97, 98, 99, 100], our RL framework can be used for correcting coherent physical errors in the code as well. Moreover, we can extend our action space by introducing additional error removal operations that can correct higher-weight errors. It is also possible to apply our RL framework to different codes not discussed in this work as well. For example, it would be interesting to consider higher-dimensional color codes, which support transversal non-Clifford logical operations [101, 102, 103], have interesting single-shot error correction properties [104, 105], and can also be realized in state-of-the-art quantum platforms [6, 106].
Although we employed a simple neural network to optimize our LEC circuit, it is also possible to use a more complicated toolkit — including transformer [107] that has already been used in designing an efficient global decoder for 2D surface code [22, 26]. The advantage of using these more advanced schemes is that we may not need to provide a fixed gate set — depending on our knowledge of error patterns on a given error-correcting code — as input. We can allow RL to find the proper gate input set from scratch by only providing the type of available gates as input ingredients.
@@ Acknowledgements.
Appendix A Circuit-level simulation of QEC
@@
  • We represent data qubits as three-dimensional arrays: first dimension for a type of a Pauli error on the qubit (either 00 for $\hat{I}$, 01 for $\hat{X}$, 10 for $\hat{Z}$, or 11 for $\hat{Y}$), second dimension for an index of a copy of a code, and third dimension for an index of a qubit within a code.
  • To perform a parallelized CNOT, CCX, or CCZ gate operation, we first define lists of qubit indices that match with the gate connectivity. For example, for parallelized CCX gates, we define two lists for ancilla indices and a list for data qubit indices. Then, the $i$-th index for each list corresponds to control and target qubits for the $i$-th CCX gate to be applied. Although we implement the Pauli error propagation for each CNOT gate, we model each ancilla classically, i.e. either |$\ket{0}$ or |$\ket{1}$, to avoid dealing with error propagation through non-Clifford CCZ and CCX gates. Physically, this condition can be approximately achieved by resetting ancillas regularly as illustrated in Fig. 1.
  • When ambient or gate errors with error rate $p_{err}$ are introduced to the qubits, we generate other arrays with the same shapes where each entry is 0 with probability $1-p_{err}$ and 1 with probability $p_{err}$ and then perform XOR operation between the original arrays and such “error arrays.” Note that for each copy of code, different patterns of errors are introduced to data qubits and ancillas. This enables us to benchmark a decoding success rate of the given LEC circuit with given error rates.
@@
However, for 4D toric code, we effectively represent the data qubits and ancillas using odd and even coordinates [108]. Let’s consider a coordinate $(x_{0},x_{1},x_{2},x_{3})$ on 4D lattice, where each $x_{i}$ has an integer value between 0 and $12L-$ 1. Then, we consider data qubits — on “faces” of the 4D lattice — are located on coordinates where two of $x_{i}$s are even and two of them are odd. Similarly, we consider ancillas that store eigenvalues of edge-type (or cube-type) stabilizers with six $\hat{Z}$ (or $\hat{X}$) operators to be located on coordinates where three of $x_{i}$s are even (or odd) and one of them is odd (or even). We denote odd coordinates as $o$ and even coordinates as $e$ to classify coordinates:
@@
This representation provides an intuitive understanding of which data qubits support each stabilizer. For example, an edge-type stabilizer on $(1,2,2,2)$ is made of $\hat{Z}$ operators on data qubits $(1,1,2,2)$, $(1,2,1,2)$, $(1,2,2,1)$, $(1,3,2,2)$, $(1,2,3,2)$, and $(1,2,2,3)$. Also, such structure of connectivity provides us an understanding of logical $\hat{Z}$ and $\hat{X}$ operators. For example, a topologically non-trivial “sheet” of $\hat{Z}$ operators on all data qubits with coordinates $(1,1,2k_{1},2k_{2})$ with $10\leq k_{1},k_{2}\leq L-$ 1 correspond to logical $\hat{Z}$ operator, which indeed commutes with all stabilizers. From this analysis, six independent logical $\hat{Z}$ and $\hat{X}$ operators can be easily identified.
@@
Based on this representation, we design a parallelized error removal operation based on CCX and CCZ gates for 4D toric code. As shown in Table 3, there are six different configurations of parallelized Toom Sweep operations, where each consists of $L^{2}$ CCX gates and $L^{2}$ CCZ gates. Since each configuration has four choices of “direction” just as in the 2D Ising model illustrated in Fig. 5(b), we have $6\times 4=24$ parallelized Toom Sweep operations. On the other hand, CCX/CCZ gates for both $d=1$ and $d=(1,1)$ operations are controlled by the same parity classification of ancillas. As shown in Table 4, there are four different configurations of parallelized $d=1$ and $d=(1,1)$ operations, where each consists of $23L^{2}/$ 2 CCX gates and $23L^{2}/$ 2 CCZ gates — this is why we choose $L$ to be even. For $d=1$ operation, each configuration has two choices for a group of data qubits. For example, $d=1$ CCX targeted to $(1,1,2,2)$ and controlled by $(1,0,2,2)$ and $(1,2,2,2)$ cannot be done in parallel with $d=1$ CCX targeted to $(1,3,2,2)$ and controlled by $(1,2,2,2)$ and $(1,4,2,2)$. Thus, we have $4\times 2=8$ parallelized $d=1$ operations. For $d=(1,1)$ operation, there are four additional degrees of freedom on top of two choices for a group of data qubits. For example, $d=1$ CCX targeted to $(1,1,2,2)$ can be controlled by $(1,0,2,2)$ and either $(1,2,0,2)$, $(1,2,4,2)$, $(1,2,2,0)$, or $(1,2,2,4)$. Thus, we have $4\times 2\times 4=32$ parallelized $d=(1,1)$ operations. In summary, we include twenty-four Toom Sweep operations, eight $d=1$ operations, and thirty-two $d=(1,1)$ operations as the candidates of error removal operations. For simplifying the RL training and reducing the number of syndrome extraction operations, we “grouped” parallelized $d=1$ and $d=(1,1)$ operations each into single actions.
@@
From such design of error removal operations, we optimize the order of applying parallelized CNOT gates for syndrome extraction operation. For 2D toric code, there are four “steps” of parallelized CNOT gates in sequence to perform a syndrome extraction as illustrated in Fig. 3. For the 2D Ising model and 4D toric code, we perform a syndrome excitation before each error removal operation. We only update syndromes for the ancillas that are used in the following error removal operation. For example, for Type 1 Toom operation in Table 3, we only prepare stabilizer ancillas on $(o,e,e,e)$, $(e,o,e,e)$, $(e,o,o,o)$, $(o,e,o,o)$, i.e. $4L^{2}$ ancillas instead of $8L^{2}$. We decompose such syndrome extraction for $4L^{2}$ ancillas into six “steps” of parallelized CNOT gates in sequence. After finding the optimal $N$ number of steps of parallelized CNOT gates in sequence, we test $N!$ possible permutations of this sequence and choose the one with the smallest average number of errors introduced during the entire syndrome extraction operation. Note that it is also possible to optimize such sequence more systematically as discussed in previous works [109, 110].
Appendix B Details on RL training
B.1 Conditions on training termination
Each epoch of our RL training consists of 500 episodes with a mini-batch size of 50. In other words, we collect 500 samples of LEC circuits and their benchmarked decoding success rates, divide them into smaller mini-batch subsets with 50 samples each, and perform 10 updates with these subsets on the policy and value network. Before saving our updated neural network model, we run 40 epochs for the 2D toric code, 80 epochs for the 2D Ising model, and 40 epochs for the 4D toric code. We terminate each RL training when a neural network model produces the same LEC circuit as output as the previously saved one throughout these epochs. For 2D toric code with $H_{\max}=40$ and $p_{gate}=1\times 10^{-4}$, each RL training requires 200-400 epochs. For self-correcting memories with $H=60$ and any $p_{gate}$, each RL training requires 800-1,000 epochs for the 2D Ising model and 400-600 epochs for the 4D toric code. Note that these are obtained with a fully connected neural network with two hidden layers of size 128. Also, we use a variable depth agent for 2D toric code and a fixed depth agent for 2D Ising model and 4D toric code as discussed in detail in Appendix B.2.
B.2 Details on agent: fixed vs. variable depth
As discussed in Section IV.2, we construct two different types of RL agents: (1) fixed-depth agent for 2D Ising model and 4D toric code and (2) variable-depth agent for 2D toric code. Note that both types are based on the scheme that given a circuit depth $H$, the RL agent builds a sequence of $H$ actions as illustrated in Fig. 6(a). This scheme is summarized in Algorithm 2, which defines the “BuildLEC” function in Algorithm 1.
@@ B.1 Conditions on training termination
A neural network $f^{\vec{\theta}}$ in the agent takes all previous actions as an input and outputs the next action. To represent the previous actions, we one-hot encoded the LEC action sequence into $|\mathcal{A}|$-by-$H$ matrix with entries 00 or $1$. For example, the LEC action sequence $\{A_{1},\ldots,A_{h}\}$ can be written by the $|\mathcal{A}|$-by-$H$ matrix with $(A_{h},h)$ entries are $1$ and all other entires are 00. Note that for 2D toric code, each action $A_{i}$ is either syndrome extraction or one of error removal operations in Fig. 5(a). Additionally, we fix that the first action $A_{1}$ is always a syndrome extraction operation for 2D toric code. For the 2D Ising model and 4D toric code, each action $A_{i}$ is one of the error removal operations in Fig. 5(b) followed by syndrome extraction operation on corresponding control ancillas. We do not fix the first action for the 2D Ising model and 4D toric code, because a syndrome extraction operation is performed for any action before an error removal operation. We optimize and fix the order of CNOT gates that consist of each syndrome extraction operation as discussed in Appendix A.
@@
Two limitations of this modified RL framework are that (1) $H_{opt}\leq H_{\max}$ and (2) $H_{opt}$ found by RL training is suboptimal if true $H_{opt}$ is much smaller than $H_{\max}$. We set initial $H_{\max}$ to be 40 for 2D toric code as we confirmed from training with the fixed-depth agent that $H_{opt}$ is smaller than 40. Then, after the termination condition is satisfied and $H_{opt}$ is obtained, we set the next $H_{\max}$ to be the $H_{opt}$ of the previous training and run the next training. This is why we emphasize in Section IV.2 that the training with a variable-depth agent is more “expensive” than the training with a fixed-depth agent.
@@
We have tried an alternative approach to optimize the LEC circuit with a variable depth by adding a stop action — instead of the null action $\emptyset$. Instead of “doing nothing,” this action terminates the LEC circuit immediately. This approach was not successful, because the RL training resulted in building the NN LEC circuit with circuit depth 5 [46], even for training $p_{gate}=1\times 10^{-4}$ where a true $H_{opt}>30$. Since the NN LEC circuit is good enough to possibly form a “deep” local minimum in the reward landscape, we believe that a stop action is incentivized enough to be called early. Although it could be possible to get around this issue by modifying the reward structure, we find that the $\emptyset$-action option is more natural to achieve the variable-depth LEC circuit optimization.
B.3 Details on RL hyperparameters
2D Toric Code 2D Ising Model
$(L,H_{\max}N,R,p_{amb},p_{gate})=(8,40,100,7,0.02,1\times 10^{-4})$ $(L,H_{\max}N,p_{amb},p_{gate})=(8,60,100,0.40,1\times 10^{-3})$
$d_{NN}$ Run 1 Run 2 Run 3 Run 4 $d_{NN}$ Run 1 Run 2 Run 3 Run 4
Table 5: RL training results for 2D toric code and 2D Ising model with various network size $d_{NN}$. We construct the RL agent with a fully connected neural network with two hidden layers of size $d_{NN}$ (for both the policy and value network of a PPO algorithm) and the RL environment with the parameters shown in the table. In each cell, we record a final reward and the required number of epochs before training termination discussed in Appendix B.1 — in the format as “Reward (# epochs)” — for each $d_{NN}$ and training run index.
As shown in Table 5, we compare the final reward from RL training with a fully connected neural network with two hidden layers of size $d_{NN}$ with four independent runs. We observe that the final reward gets improved until $d_{NN}=128$. However, for $d_{NN}>128$, we observe that a similar reward was obtained with shorter training epochs before termination. Considering the size of saved neural networks after training, we choose $d_{NN}=128$ for our RL training. Also, this is why we did not have to implement a more complicated neural network structure — such as a transformer — because a simple fully connected neural network with two hidden layers of size 128 was enough for our optimization task.
B.4 Details on environment
As illustrated in Fig. 6 and described further in Section IV.1, we simulate $R$ number of LEC rounds — introduction of ambient errors followed by LEC circuits with gate errors — for $N$ times in parallel before computing the reward. For both reward and lifetime computation, we determine the decoding success/failure followed by the final recovery step as below:
  • 2D toric code: Based on the precise eigenvalues of all stabilizers, perform a minimum-weight perfect matching (MWPM) classical algorithm to remove all remaining syndromes and decode the final state to the ground state. We used the Python-based package PyMatching to perform this algorithm in our implementation of RL framework [14, 16]. If we get a trivial ground state, where all remaining data qubit errors can be written as a product of stabilizers, the decoding by LEC is successful. If we get a ground state with logical error(s), which are non-trivial loop(s) of data qubit errors, the decoding by LEC fails.
  • 2D Ising model: Perform a majority vote among data qubits. If less than 50% of them are bit-flipped, the decoding by LEC is successful. Otherwise, the decoding by LEC is failed.
  • 4D toric code: Perform enough (50 rounds) repetition of perfect Toom’s rule decoding. If we get a trivial ground state, where all remaining data qubit errors can be written as a product of stabilizers, the decoding by LEC is successful. If we fail to get a ground state (due to uncorrectable errors by Toom’s rule LEC) or we get a ground state with logical error(s), which are non-trivial sheet(s) of data qubit errors, the decoding by LEC is failed.
As discussed in Section III.2, the uncorrectable error configurations after perfect Toom’s rule decoding in 4D toric code can have a low error density. Thus, in principle, these uncorrectable errors should be decoded by some non-local decoding protocol, which is discussed in [44] for example. However, we decide to leave these uncorrectable errors as decoding failures, because (1) the non-local decoding protocol takes significant running time and (2) the prior LEC circuits prevent the uncorrectable error configurations with low error density.
For 2D and 4D toric code, the reward is computed as the number of successful decoding cases divided by the number of total cases $N$. However, for the 2D Ising model, the reward is computed as an average non-error density, i.e. average number of non-flipped spins divided by the total number of spins $L^{2}$, after repeating $R$ LEC rounds. This reward computation method is summarized in Algorithm 3.

### v2

@@ VI.1 Reducing rate of mid-circuit readouts
Section V shows that LEC circuits can be used to remove errors from encoded qubits and to extend the memory lifetime. Then, it is reasonable to ask whether such LEC circuits can be beneficial to extend memory lifetime in practice. We focus on a 2D toric code to investigate the practical application. In 2D toric code, using only LEC has limitations for a practical application, because memory lifetime extended by LEC does not scale with $L$ as observed in Fig. 9. Also, undecoded error chains get accumulated by the repetition of LECs, which implies that extending memory lifetime has some upper limit. However, we can envision a decoding strategy combining the advantages of both LEC and global decoding. In particular, repeated application of LEC can be used to reduce the rate of error accumulation of an encoded logical qubit. Then, we can periodically perform the global decoding. The key advantage of this approach is that LEC circuits can be applied more quickly compared to the global decoding that involves mid-circuit readouts.
Figure 11: Original vs. Hybrid decoding protocol to preserve 2D toric code memory. (a) At each unit time, new errors with rate $p_{\rm{unit}}=0.1\%$ are introduced to data qubits. Throughout total physical time $T$, global decoding (blue, solid lines) is performed regularly for both original (blue circle) and hybrid protocol (orange diamond). However, hybrid decoding protocol regularly applies LEC circuit (orange, dotted lines) between global decoding applications. After $T$ passes, we evaluate the final logical error rate $P_{\rm{L}}$ (see Appendix D for the error bar). (b) Minimum number of global decoding vs. $p_{\rm{unit}}\times T$ in each protocol for given multi-qubit gate fidelities and $P_{\rm{L}}$ (see Appendix F for details).
@@
In practice, we aim for the memory to exhibit a logical error rate $P_{\rm{L}}$ below some target value after a given physical time $T$ where each unit time step introduces some error with rate $p_{\rm{unit}}$ to data qubits. We compare original and hybrid decoding protocols as shown in Fig. 11(a). The original decoding protocol performs $N_{\rm{Global}}^{(\rm{Original})}$ global decoding with equally spaced time intervals throughout the total time $T$. The hybrid decoding protocol performs $N_{\rm{Global}}^{(\rm{Hybrid})}$ global decoding with RL-optimized LEC circuits interleaved between global decoding. Our goal is to find the regime of parameters where
$N_{\rm{Global}}^{\rm{(Original)}}>N_{\rm{Global}}^{\rm{(Hybrid)}}.$ (8)
@@
When two-qubit gate fidelity $F_{\rm{2Q}}=99.96\%$, three-qubit gate fidelity $F_{\rm{3Q}}=99.88\%$, and target $P_{\rm{L}}<15\%$, two protocols contrast as shown in Fig. 11(b). For $L=8$ lattice, the hybrid decoding can reduce the number of global decoding for the constant amount regardless of $T$. In other words, the gain of using the hybrid decoding protocol gets decreased as we want to preserve the memory for a longer time. However, for $L=16$ lattice, we observe that the hybrid decoding can reduce the number of global decoding by about half regardless of how long we want to preserve the memory. Note that each global decoding of 2D toric code requires $L$ repetitions of mid-circuit readout of ancillas for faulty syndrome extraction before the 3D minimum-weight perfect matching algorithm [9, 10, 16]. Thus, we can conclude that using an LEC circuit reduces the number of mid-circuit readouts by $L\times(N_{\rm{Global}}^{\rm{(Original)}}-N_{\rm{Global}}^{\rm{(Hybrid)}})$ throughout the entire physical time, i.e. we can replace some portion of the mid-circuit readouts into possibly more cost-efficient operations. Our investigation confirms that there exists a regime of high multi-qubit gate fidelity where LEC can be practically useful for a memory lifetime enhancement.
VI.2 Dissipatively preparing topological phases
@@ VII Discussion and Outlook
In this work, we develop an RL framework to optimize measurement-free error-correcting circuits, composed of a finite number of faulty multi-qubit gates. We specifically focus on a finite-sized 2D toric code, 2D Ising model, and 4D toric code. By introducing a notion of effective code distance $D_{\rm{eff}}$, we find that when gate fidelities are high enough, the RL-optimized LEC circuits are better than conventional LEC circuits in extending the lifetime of a memory encoded in these codes. Then, we additionally investigate two potential applications of this LEC scheme. First, we show that the portion of global decoding, which consists of mid-circuit readout and classical decoding, can be replaced with LEC circuits with high-fidelity multi-qubit gates for preserving 2D toric code memory. Also, we discuss that quantum states with topological phases can be prepared dissipatively.
@@
Our RL framework is adaptive to the needs of various changes. For example, keeping the same gate operations as an input action space, we can simply substitute a biased, realistic error model into our current noise channel. Since coherent errors turn into Pauli errors when we trace out the ancillas at the ancilla reset step [100, 101, 102, 103], our RL framework can be used for correcting coherent physical errors in the code as well. Moreover, we can extend our action space by introducing additional error removal operations that can correct higher-weight errors. It is also possible to apply our RL framework to different codes not discussed in this work as well. For example, it would be interesting to consider higher-dimensional color codes, which support transversal non-Clifford logical operations [104, 105, 106], have interesting single-shot error correction properties [107, 108], and can also be realized in state-of-the-art quantum platforms [6, 109].
Although our RL framework benefits from a simple and small neural network in terms of the training cost, we can choose a more complicated optimizer — including transformer [110] that has already been used in designing an efficient global decoder for 2D surface code [22, 26]. The advantage of using these more advanced schemes is that we may not need to provide a fixed gate set — depending on our knowledge of error patterns on a given error-correcting code — as input. We can allow RL to find the proper gate input set from scratch by only providing the type of available gates as input ingredients. This future direction can also be useful to obtain a general LEC circuit optimizer with the stabilizer structure of the code as an input to the RL agent such that we do not have to re-train the agent for each code.
@@ Appendix A Experimental implementation of LEC
Appendix B Circuit-level simulation of QEC
@@
  • We represent data qubits as three-dimensional arrays: first dimension for a type of a Pauli error on the qubit, second dimension for an index of a copy of a code, and third dimension for an index of a qubit within a code.
  • To perform a parallelized CNOT, CCX, or CCZ gate operation, we define lists of qubit indices that match with the gate connectivity. The $i$-th index for each list corresponds to control or target qubits for the $i$-th multi-qubit gate to be applied. Although we implement the Pauli error propagation for each CNOT gate, we model each ancilla classically, i.e. either |$\ket{0}$ or |$\ket{1}$, to avoid dealing with error propagation through non-Clifford CCZ and CCX gates. Physically, this condition can be approximately achieved by resetting ancillas regularly as illustrated in Fig. 1.
@@
However, for 4D toric code, we effectively represent the data qubits and ancillas using odd and even coordinates [114]. Let’s consider a coordinate $(x_{0},x_{1},x_{2},x_{3})$ on 4D lattice, where each $x_{i}$ has an integer value between 0 and $12L-$ 1. We denote odd coordinates as $o$ and even coordinates as $e$ to classify qubits:
@@
From this representation, we can easily identify six independent logical $\hat{Z}$ and $\hat{X}$ operators. For example, a topologically non-trivial “sheet” of $\hat{Z}$ operators on all data qubits with coordinates $(1,1,2k_{1},2k_{2})$ with $10\leq k_{1},k_{2}\leq L-$ 1 correspond to logical $\hat{Z}$ operator, which indeed commutes with all stabilizers.
@@
Now, we introduce error removal operations based on CCX and CCZ gates in parallel for 4D toric code. As shown in Table 2, there are six different configurations of parallelized Toom Sweep operations, where each consists of $L^{2}$ CCX gates and $L^{2}$ CCZ gates. Since each configuration has four choices of “direction” just as in the 2D Ising model illustrated in Fig. 5(b), we have $6\times 4=24$ parallelized Toom Sweep operations. On the other hand, CCX/CCZ gates for both $d=1$ and $d=(1,1)$ operations are controlled by the same parity classification of ancillas. As shown in Table 3, there are four different configurations of parallelized $d=1$ and $d=(1,1)$ operations, where each consists of $23L^{2}/$ 2 CCX gates and $23L^{2}/$ 2 CCZ gates — this is why we choose $L$ to be even. For $d=1$ operation, each configuration has two choices for a group of data qubits. Thus, we have $4\times 2=8$ parallelized $d=1$ operations. For $d=(1,1)$ operation, there are four additional degrees of freedom on top of two choices for a group of data qubits. Thus, we have $4\times 2\times 4=32$ parallelized $d=(1,1)$ operations. In summary, we include twenty-four Toom Sweep operations, eight $d=1$ operations, and thirty-two $d=(1,1)$ operations as the candidates of error removal operations. For simplifying the RL training and reducing the number of syndrome extraction operations, we “grouped” parallelized $d=1$ and $d=(1,1)$ operations each into single actions.
@@
From such design of error removal operations, we optimize the order of applying parallelized CNOT gates for syndrome extraction operation. For 2D toric code, there are four “steps” of parallelized CNOT gates in sequence to perform a syndrome extraction as illustrated in Fig. 3. For the 2D Ising model and 4D toric code, we perform a syndrome excitation before each error removal operation. We only update syndromes for the ancillas that would be used in the following error removal operation. After finding the optimal $N$ number of steps of parallelized CNOT gates in sequence, we test $N!$ possible permutations of this sequence and choose the one with the smallest average number of errors introduced during the entire syndrome extraction operation. Note that it is also possible to optimize such sequence more systematically as discussed in previous works [115, 116].
Appendix C Details on RL training
C.1 Details on agent
We summarize the RL agent’s construction of an LEC circuit as in Algorithm 2, which defines the “BuildLEC” function in Algorithm 1.
@@ C.1 Details on agent
The size of the action space $|\mathcal{A}|$ depends on the code: $|\mathcal{A}|=24$ for 2D toric code, $|\mathcal{A}|=12$ for 2D Ising model, and $|\mathcal{A}|=26$ for 4D toric code (see Appendix B). Also, in practice, we use two fully connected hidden layers with 128 nodes each for both policy and value networks. We provide a justification of not using bigger networks in Appendix C.4.
To represent the previous actions as an observation for the RL agent, we one-hot encoded the LEC action sequence into $|\mathcal{A}|$-by-$H$ matrix with entries 00 or $1$. For 2D toric code, each action $A_{i}$ is either syndrome extraction or one of error removal operations in Fig. 5(a), and we fix that the first action $A_{1}$ is always a syndrome extraction operation. For the 2D Ising model and 4D toric code, each action $A_{i}$ is one of the error removal operations in Fig. 5(b) followed by syndrome extraction operation on corresponding ancillas. We do not fix the first action for the 2D Ising model and 4D toric code, because a syndrome extraction operation is performed for any action before an error removal operation.
@@
Two limitations of this modified RL framework are that (1) $H_{\rm{opt}}\leq H_{\max}$ and (2) $H_{\rm{opt}}$ found by RL training is suboptimal if true $H_{\rm{opt}}$ is much smaller than $H_{\max}$. We set initial $H_{\max}$ to be 40 for 2D toric code as we confirmed from training with the fixed-depth agent that $H_{\rm{opt}}$ is smaller than 40. Then, after the termination condition is satisfied and $H_{\rm{opt}}$ is obtained, we set the next $H_{\max}$ to be the $H_{\rm{opt}}$ of the previous training and run the next training. This is why we emphasize in Section IV.2 that the training with a variable-depth agent is more “expensive” than the training with a fixed-depth agent.
@@
We have tried an alternative approach to optimize the LEC circuit with a variable depth by adding a stop action that terminates the LEC circuit immediately — instead of the $\emptyset$ action. This approach was not successful, because the RL training resulted in building the NN LEC circuit with circuit depth 5 [46], even for training $p_{\rm{gate}}=1\times 10^{-4}$ where a true $H_{\rm{opt}}>30$. Since the NN LEC circuit is good enough to possibly form a “deep” local minimum in the reward landscape, we believe that a stop action is incentivized enough to be called early. Although it could be possible to get around this issue by modifying the reward structure, we find that the $\emptyset$-action option is natural to achieve the variable-depth LEC circuit optimization.
C.2 Details on environment
We summarize the RL environment’s computation of reward as in Algorithm 3, which defines the “ComputeReward” function in Algorithm 1.
[…]
$L$ $p_{\rm{amb}}$ $N$ $R$
2D toric code 8 0.02 100 5
2D Ising model 8 0.40 100 1
4D toric code 4 0.03 50 2
Table 4: Code and error model parameters used for optimizing the LEC circuit with our RL framework. $L$ denotes the system size, $p_{\rm{amb}}$ denotes the ambient error rate in each LEC cycle before applying the LEC circuit, $N$ denotes the number of copies of code, and $R$ denotes the number of repeated LEC cycles used to compute decoding success rate as illustrated in Fig. 6.